# Simultaneously search for multi-target Galactic binary gravitational waves in reduced parameter space with LMPSO-CV

### v1

@@ Appendix A The case study of the method
We employed our method to analyze specific sources and observed its effect. In the Tab. 2, we present the parameter of binary $A$, which is the same as Fig. 6.
$f$ $\dot{f}$ $\beta$ $\lambda$ $\mathcal{A}$ $\iota$ $\psi$ $\phi_{0}$
Binary $A$ 0.002090825 $5.798272\times 10^{-17}$ 0.034327 4.809826 $3.127627\times 10^{-22}$ 1.036248 0.228909 3.22735530717959
@@
For the test, we initially generate the gravitational waveform of binary $A$ in the LISA data and then introduce instrument noise to create mock data. Since there is only one gravitational wave signal in the data, the task of identifying the signal and removing noise is straightforward. In Fig. 21, we present the parameter positions of the injected source (binary $A$) and the reported source identified by our method in normalized parameter space. It is evident that the positions of the two sources coincide, confirming the accuracy and effectiveness of our search. The correlation coefficient for the reported source is 0.9998.
Figure 21: The red cross signifies the position of binary $A$ in the parameter space, while the blue dot represents the reported source identified by our method. The ellipsoid depicts the void with the reported source as the center.
We also assess the performance of our method in scenarios where two signals overlap. We generate gravitational waveforms for binaries $A$ and $B$ in the LISA detector, with the frequency of binary $B$ incremented by $4\times\mathrm{d}f$ relative to binary $A$. The resulting waveforms, combined with instrument noise, constitute the mock data, corresponding to Fig. 7(c) and Fig. 7(d). Given the overlap of the signals, resolving the two sources becomes more challenging. Due to the low correlation between the two search results during cross-validation, we set the $R_{ee}$ threshold to 0.7. Ultimately, our method identifies four reported sources, comprising one accurate source, one contaminated source, and two instances of unremoved degeneracy noise. The correlation coefficients for these reported sources are 0.9139, 0.6010, 0.4976, and 0.3510, respectively.
@@
Figure 22: The red cross on the left represents binary $A$, while the cross on the right represents binary $B$. The four blue dots represent the reported sources identified by our method. The ellipsoids represent the voids with these reported sources as their centers.

### v3

@@ Appendix A The case study of the method
We employed our method to analyze specific sources and observed its effects. In Tab. 3, we present the parameters of binary $A$, which were used to generate Fig. 5 and Fig. 6. For the test, we first generated the gravitational waveform of binary $A$ in the LISA data, then introduced instrument noise to create mock data. Since there is only one gravitational wave signal in the data, identifying the signal and removing noise is straightforward. In Fig. 20(a), we present the parameter positions of the injection source (binary $A$) and the source identified by our method in normalized parameter space. It is evident that the positions of the two sources coincide, confirming the accuracy and effectiveness of our search. The correlation coefficient for the identified source is 0.9998, as shown in Tab. 4.
$f$ $\dot{f}$ $\beta$ $\lambda$ $\mathcal{A}$ $\iota$ $\psi$ $\phi_{0}$ SNR
Binary $A$ 0.002090825 $5.798272\times 10^{-17}$ 0.034327 4.809826 $3.127627\times 10^{-22}$ 1.036248 0.228909 3.227355 139.3
@@
We also evaluate our method’s performance in scenarios involving overlapping signals. Gravitational waveforms are generated for binaries $A$ and $B$ in the LISA detector, with the frequency of binary $B$ increased by $4\times\mathrm{d}f$ relative to binary $A$, corresponding to the configurations in Fig. 7(c) and Fig. 7(d). These waveforms, combined with instrument noise, form the mock data. Fig. 21 illustrates the $\mathcal{F}$-statistic distribution in the three-dimensional parameter space for this case, where the two crosses represent the parameter locations of binaries $A$ and $B$. In the first scenario, we searched for all local maxima that met the specified criteria. This search identified two reported sources with nearly identical frequencies, corresponding to two local maxima in subgraph (3, 2) of Fig. 21. Both of these maxima had $\mathcal{F}$-statistic values exceeding those at the true parameter positions, with correlation coefficients of 0.3976 and 0.4128. In the second scenario, these two local maxima were excluded from the search results, yielding two reported sources directly aligned with the true parameter positions and correlation coefficients of 0.9240 and 0.9266, as shown in Tab. 4. For both scenarios, we present the parameter positions of the injection sources (binaries $A$ and $B$) and those identified by our method in the normalized parameter space in Fig. 20(b) and Fig. 20(c), respectively.
We generated gravitational waveforms for binaries $A$ and $B$, with the frequency of binary $B$ increased by $5\times\mathrm{d}f$ relative to binary $A$. Fig. 22 shows the distribution of the $\mathcal{F}$-statistic in three-dimensional parameter space for the two sources, which is corresponding to the situation in Fig. 7(a) and Fig. 7(b). Since no local maximum in the $\mathcal{F}$-statistic exceeds the value at the true positions of the two injection sources, the $-\mathcal{F}$-statistic analysis proceeds smoothly, and the two sources are successfully identified. In Fig. 20(d), we present the parameter positions of the injection sources (binaries $A$ and $B$) and the sources identified by our method in normalized parameter space. The correlation coefficients for the two reported sources are 0.9847 and 0.9927, respectively, as shown in Tab. 4.
$f$ $\dot{f}$ $\beta$ $\lambda$ $\mathcal{A}$ $\iota$ $\psi$ $\phi_{0}$ SNR $R$
Fig. 20(a) 0.002090825 $6.345725\times 10^{-17}$ 0.037617 4.810014 $3.041765\times 10^{-22}$ 1.011543 0.249035 3.182173 139.5 0.9998
Fig. 20(b) 0.002090858 $4.438414\times 10^{-17}$ 0.028935 4.508666 $2.389844\times 10^{-22}$ 1.027265 0.271220 4.485963 107.8 0.3976
0.002090857 $7.105742\times 10^{-17}$ 0.006471 5.117904 $2.474984\times 10^{-22}$ 1.036071 0.587262 0.837439 111.6 0.4128
Fig. 20(c) 0.002090825 $4.489096\times 10^{-17}$0.255309 4.827474 $2.441402\times 10^{-22}$ 1.127144 0.610877 2.434201 98.9 0.9240
0.002090889 $1.027816\times 10^{-16}$ 0.303494 4.814666 $2.143366\times 10^{-22}$ 1.011473 0.662798 2.245402 98.7 0.9266
Fig. 20(d) 0.002090829 $-2.099675\times 10^{-17}$ 0.058268 4.805761 $3.129668\times 10^{-22}$ 1.031005 0.219305 2.782279 140.4 0.9847
0.002090904 $2.530915\times 10^{-17}$ 0.028884 4.809350 $3.180976\times 10^{-22}$ 1.051616 0.222846 3.364739 139.5 0.9927
Table 4: The parameters of reported sources identified by our method in the case study.
@@
Figure 20: The red crosses signify the position of injection sources in the parameter space, while the blue dots represent the reported sources identified by our method. The ellipsoids depict the voids with the reported sources as the center.
Figure 21: Slices of the $\mathcal{F}$-statistic in three-dimensional parameter space for binary $A$ and $B$. The parameter position of binary $A$ is marked in the subfigure (2, 5) with a red cross, and the parameter position of binary $B$ is marked in the subfigure (3, 4) with a red cross. The frequency intervals between two adjacent subfigures correspond to the minimum frequency resolution $\mathrm{d}f=1.59\times 10^{-8}$ Hz. In each subfigure, the horizontal coordinate represents Ecliptic Longitude $\lambda$, and the vertical coordinate is Ecliptic Latitude $\beta$. The figure does not encompass the entire frequency range of the degeneracy noise distribution.
Figure 22: Slices of the $\mathcal{F}$-statistic in three-dimensional parameter space for binary $A$ and $B$. The parameter position of binary $A$ is marked in the subfigure (2, 5) with a red cross, and the parameter position of binary $B$ is marked in the subfigure (3, 5) with a red cross. The frequency intervals between two adjacent subfigures correspond to the minimum frequency resolution $\mathrm{d}f=1.59\times 10^{-8}$ Hz. In each subfigure, the horizontal coordinate represents Ecliptic Longitude $\lambda$, and the vertical coordinate is Ecliptic Latitude $\beta$. The figure does not encompass the entire frequency range of the degeneracy noise distribution.